# CtrSVDD: A Benchmark Dataset and Baseline Analysis for Controlled Singing Voice Deepfake Detection

### v1

@@ Abstract
Recent singing voice synthesis and conversion advancements necessitate robust singing voice deepfake detection (SVDD) models. Current SVDD datasets face challenges due to limited controllability, diversity in deepfake methods, and licensing restrictions. Addressing these gaps, we introduce CtrSVDD, a large-scale, diverse collection of bonafide and deepfake singing vocals. These vocals are synthesized using state-of-the-art methods from publicly accessible singing voice datasets. CtrSVDD includes 47.64 hours of bonafide and 260.34 hours of deepfake singing vocals, spanning 14 deepfake methods and involving 164 singer identities. We also present a baseline system with flexible front-end features, evaluated against a structured train/dev/eval split. The experiments show the importance of feature selection and highlight a need for generalization towards deepfake methods that deviate further from training distribution. The CtrSVDD dataset and baselines are publicly accessible [^1] https://github.com/SVDDChallenge/CtrSVDD2024_Baseline.
@@ 1 Introduction
Research has emerged towards the singing voice deepfake detection (SVDD) task as a response to synthesized singing voices. We previously proposed a multilingual in-the-wild dataset, SingFake [5], by collecting deepfake song clips from user-generated content websites. The label of bonafide or deepfake was identified by the uploaders and manually verified by the annotators. The synthesis methods utilized rely on uploaders to disclose and are often lacking: 60.6% of deepfake songs in SingFake are reported with the “Unknown” generation method. Furthermore, out of all song clips that reported the generation method in SingFake, 92.2% reported variants of SoVITS  [^2] https://github.com/svc-develop-team/so-vits-svc, indicating a potential lack of diversity. Xie et al. [6] curated a controlled Chinese song dataset for fake song detection (FSD). The deepfake songs are generated using one SVS and four SVC methods applied to the bonafide songs they collected. Due to licensing restrictions, its bonafide set is not publicly available. Both works also found that speech-trained deepfake detection models cannot directly work on the SVDD task, highlighting the need for a singing voice deepfake detection dataset.
@@
With the CtrSVDD dataset, we also present baseline systems for CtrSVDD with flexible front-end modules (encoding waveforms into feature representations) and a fixed, robust back-end module (making predictions). Using this baseline, we explored the impact of front-end features by comparing raw waveform, spectrogram-based, and cepstral coefficients (CC)-based features. The CtrSVDD dataset, baseline system implementations, and trained model weights are publicly accessible, and their links will be provided in the camera-ready version.
@@ 2.2 Details of deepfake generation methods
Figure 1: Histogram of audio duration. The left subfigure shows the distribution across two classes (bonafide and deepfake), whereas the right one differentiates among the train/dev/eval splits. We exclude data exceeding three standard deviations from the mean (0.6% of all data) for better visualization.
@@ 4.1 Experimental setup
In evaluation, we applied 5 different random seeds to trim vocal clips, creating 5 variations of the test set. We report the mean and standard deviation of the EER across these versions to assess model robustness against random time shifts. All experiments were performed on a single RTX 4090 GPU. The training time for the raw-waveform-based model is slightly longer than 24 hours, while the spectrogram-based model trains for around 7 hours and all other frontend features for about 5 hours.

### v2

@@ Abstract
Recent singing voice synthesis and conversion advancements necessitate robust singing voice deepfake detection (SVDD) models. Current SVDD datasets face challenges due to limited controllability, diversity in deepfake methods, and licensing restrictions. Addressing these gaps, we introduce CtrSVDD, a large-scale, diverse collection of bonafide and deepfake singing vocals. These vocals are synthesized using state-of-the-art methods from publicly accessible singing voice datasets. CtrSVDD includes 47.64 hours of bonafide and 260.34 hours of deepfake singing vocals, spanning 14 deepfake methods and involving 164 singer identities. We also present a baseline system with flexible front-end features, evaluated against a structured train/dev/eval split. The experiments show the importance of feature selection and highlight a need for generalization towards deepfake methods that deviate further from training distribution. The CtrSVDD dataset [^1] https://zenodo.org/records/10467648 (Train/Dev) [^2] https://zenodo.org/records/10742049 (Eval) and baselines [^3] https://github.com/SVDDChallenge/CtrSVDD2024_Baseline are publicly accessible.
@@ 1 Introduction
Research has emerged towards the singing voice deepfake detection (SVDD) task as a response to synthesized singing voices. Our previous work introduced a multilingual in-the-wild dataset, SingFake [5], by collecting deepfake song clips from user-generated content websites. The label of bonafide or deepfake was identified by the uploaders and manually verified by the annotators. The synthesis methods utilized rely on uploaders to disclose and are often lacking: 60.6% of deepfake songs in SingFake are reported with the “Unknown” generation method. Furthermore, out of all song clips that reported the generation method in SingFake, 92.2% reported variants of SoVITS  [^4] https://github.com/svc-develop-team/so-vits-svc, indicating a potential lack of diversity. Xie et al. [6] curated a controlled Chinese song dataset for fake song detection (FSD). The deepfake songs are generated using one SVS and four SVC methods applied to the bonafide songs they collected. Due to licensing restrictions, its bonafide set is not publicly available. Both works found that speech-trained deepfake detection models cannot directly work on the SVDD task, highlighting the need for a singing voice deepfake detection dataset.
@@
With the CtrSVDD dataset, we also present baseline systems for CtrSVDD with flexible front-end modules (encoding waveforms into feature representations) and a fixed, robust back-end module (making predictions). Using this baseline, we explored the impact of front-end features by comparing raw waveform, spectrogram-based, and cepstral coefficients (CC)-based features. The CtrSVDD dataset, baseline system implementations, and trained model weights are publicly accessible.
@@ 2.2 Details of deepfake generation methods
Figure 1: Histogram of audio duration. The left subfigure shows the distribution across two classes (bonafide and deepfake), whereas the right one differentiates among the train/dev/eval splits. We exclude data exceeding three standard deviations from the mean (0.6% of all data) for better visualization. All distributions are visualized with a 50% opacity then overlapped for a direct comparison between them.
@@ 4.1 Experimental setup
In evaluation, we apply 5 different random seeds to trim vocal clips, creating 5 variations of the test set. We report the mean and standard deviation of the EER across these versions to assess model robustness against random time shifts. All experiments are performed on a single RTX 4090 GPU. The training time for the raw-waveform-based model is slightly longer than 24 hours, while the spectrogram-based model trains for around 7 hours and all other frontend features for about 5 hours.